A charge Q is located at a grid intersection in th
Answer #1

A charge Q is located at a grid intersection in the x-y plane which is given below as :

4,4 4.2 3.8 5.6 5.8 6 6.26.4 x (n)

The electric potential in the x-y plane due to a charge distribution near the origin is given by,

V (x,y) = \beta x / (x2 + y2)3/2

where, \beta = 5.347 x 105 Nm3/C

x = 0.4m & y = 0.4m

then, we get

V = (5.347 x 105 Nm3/C) (0.4 m) / [(0.4 m)2 + (0.4 m)2]3/2

V = (2.138 x 105 Nm4/C) / (0.32 m2)3/2

V = (2.138 x 105 Nm4/C) / (0.181 m3)

V = 1.18 x 106 V

Part-A : The x-component of the electric force which will be given as :

Fx = q E

where, E = electric field = V / d

Fx = q (V / d) = (3.3 x 10-6 C) (1.18 x 106 V) / (0.565 m)                           (d = \sqrt{}x2 + y2)

Fx = 6.89 N

Part-B : The y-component of force on Q will be given as -

using a formula, we have

W = q \DeltaV

Fy d = q (VN - VU)

Fy = (3.3 x 10-6 C) [(8.956 x 103 V) - (8.167 x 103 V)] / (0.565 m)

Fy = 4.60 x 103 N